OpenStudy (commonlizer):

What is the value of y in the equation 2 + y = −3? −1 −5 1 5

OpenStudy (commonlizer):

I think it is -1 am I correct

OpenStudy (toxicwaste):

yeah

alones (alones):

No -1 is wrong. First of all simplify both sides of the equation. y+2=−3 than subtract 2 from both sides. y+2−2=−3−2 y=−5

OpenStudy (toxicwaste):

ohh I get it

alones (alones):

Did you guess?

OpenStudy (commonlizer):

No my math is really terrible

alones (alones):

Well, hope you see the mistake now cx

OpenStudy (commonlizer):

thank you @AloneS

alones (alones):

My pleasure ;)